OpenStudy (anonymous):

Can someone please help me write out an equation to solve this question? Q. Six people are sitting in a room. When an 80 year old man enters the room, the average age of the people in the room doubles. What was the average age of the six people originally sitting in the room?

OpenStudy (anonymous):

6 + 80 = x x ÷ 2 = answer

OpenStudy (anonymous):

X= 34? So x is representing the average age of the 6 people?

OpenStudy (anonymous):

Oops 43**

OpenStudy (anonymous):

Wait let me think out loud 86 made it double so before it would have been 43 soooo 43 ÷ 6

OpenStudy (unklerhaukus):

i dont think thats right becca

OpenStudy (anonymous):

I don't think that's right actually

OpenStudy (anonymous):

Dont we have to work with fractions or something?

OpenStudy (anonymous):

let "x" be sum of ages of 6 ppl, "t" their average age. x/6 = t (x+80)/7 = 2t

OpenStudy (anonymous):

two equations and two unknowns. you can solve them... ? question is asking to find t

OpenStudy (anonymous):

Oh! Thank you @ramya25 that helped me a lot

OpenStudy (anonymous):

np :)

OpenStudy (anonymous):

Sum of ages of 6 people = 6*A Sum of ages of 6 people + X = 7*(2A) 6*A + 80 = 14A 8A = 80 A = 10 Average = 60..

OpenStudy (anonymous):

T = 10!!

OpenStudy (anonymous):

Yes Average = 10..

OpenStudy (anonymous):

Thanks everyone!

OpenStudy (anonymous):

Not 60 It is 10 I have written wrong there...

OpenStudy (anonymous):

HAHA I still understand it, so it's okay (: